OpenStudy (anonymous):

EASY EASY MEDAL Would it be the third answer? Simplify 6 to the 5th over 7 to the 3rd all raised to the 2nd power. 6 to the 7th over 7 to the 10th 6 to the 10th over 7 to the 6th 6 to the 3rd over 7 12 to the 5th over 14 to the 3rd

OpenStudy (anonymous):

\[(\frac{ 6^5 }{ 7^3 })^2=(\frac{ 7776 }{ 343 })^2=513.9540158\]

OpenStudy (anonymous):

So it would be B?

OpenStudy (lynfran):

yes

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[(\frac{ 6^5 }{ 7^3 })^2=\frac{ 6^{5\times2} }{ 7^{3\times2} }=\frac{ 6^{10} }{ 7^6 }\]